# Green’s functions, propagation invariants, reciprocity theorems, wave-field representations and propagator matrices in 2D time-dependent materials

### v1

@@ 1 Introduction
Temporal changes of material parameters have an effect on wave propagation and scattering, comparable with, but not identical to, the effect of spatial changes. Since the initial work on wave propagation and scattering in time-dependent materials (Morgenthaler, 1958; Jiang, 1975), the research in this field has recently got significant momentum (Bacot et al., 2016; Fink & Fort, 2017; Koutserimpas & Fleury, 2018; Bal et al., 2019; Peng et al., 2020; Caloz & Deck-Léger, 2020a; Mounaix et al., 2020; Ramaccia et al., 2020; Garnier, 2021; Apffel & Fort, 2022; Moussa et al., 2023; Hidalgo-Caballero et al., 2023; Ptitcyn et al., 2023). Various authors discuss the analogy between the wave equations for time-dependent and space-dependent materials (Mendonça & Shukla, 2002; Xiao et al., 2014; de Hoop & Lager, 2014; Salem & Caloz, 2015; Torrent et al., 2018; Caloz & Deck-Léger, 2020b; Van Manen et al., 2024). For 1D applications, the roles of the time- and space-coordinates are interchanged between the wave equations for both types of material. However, causality conditions apply to the time coordinate in both types of material and hence they are not interchanged. As a consequence, solutions of the wave equation in a time-dependent material are, in general, not one-to-one exchangeable with those in a space-dependent material. In a recent paper (Wapenaar et al., 2024) we systematically analyze the similarities and differences of waves in 1D space-dependent and time-dependent materials. In multi-dimensional materials, the number of space dimensions (two or three) is different from the number of time dimensions (one) and hence the analogy between the wave equations for both types of material is less straightforward. In the current paper we discuss fundamental aspects of wave propagation and scattering in 2D time-dependent materials. The discussion partly reviews work discussed in the references mentioned above but also contains new results. Like in our paper on 1D materials, we use a unified notation which simultaneously captures electromagnetic, acoustic and elastodynamic shear waves. We discuss transmission and reflection coefficients, Green’s functions, propagation invariants, reciprocity theorems, wave field representations, Green’s function retrieval, propagator matrices and the relation between the Green’s function and the propagator matrix of a time-dependent material.
@@ 2 Unified wave equation
Table 1: Specification of the quantities in equations (1) – (4). For TE (transverse electric) and TM (transverse magnetic) waves, the quantities are electric and magnetic flux densities ${D}$ and ${B}$, electric and magnetic field strengths ${E}$ and ${H}$, permittivity $\varepsilon$, permeability $\mu$, and external electric and magnetic current densities ${J}^{\rm e}$ and ${J}^{\rm m}$. For acoustic waves they are cubic dilatation $\Theta$, mechanical momentum density ${m}$, acoustic pressure $p$, particle velocity ${v}$, compressibility $\kappa$, mass density $\rho$, volume-injection rate density $q$ and external force density ${f}$. For SH (horizontally polarised shear) waves, $m$, $v$, $\rho$ and $f$ are defined the same as for acoustic waves, and the additional quantities are strain $e$, stress $\tau$, shear modulus $\mu$ and external deformation rate density $h$.
@@ 3. Acoustic $-\Theta$ $m_{x}$ $m_{z}$ $p$ $v_{x}$ $v_{z}$ $\kappa$ $\rho$ $q$ $f_{x}$ $f_{z}$
where $\alpha$ and $\beta$ are material parameters. They are also specified in Table 1 for the different wave phenomena. We can use these equations to eliminate two of the four wave field quantities from equations (1) and (2).
Let us first consider the common situation of a material which is space-dependent and constant in time, with parameters $\alpha({\bf x})$ and $\beta({\bf x})$. This implies (using equations (3) and (4)) $\partial_{t}U=\partial_{t}(\alpha P)=\alpha\partial_{t}P$ and $\partial_{t}{\bf V}=\partial_{t}(\beta{\bf Q})=\beta\partial_{t}{\bf Q}$. Hence, for this situation elimination of $U$ and ${\bf V}$ from equations (1) and (2) yields
@@
This well-known system of equations for the wave field quantities $P$ and ${\bf Q}$ underlies wave propagation in inhomogeneous, time-independent materials.
In this paper we consider homogeneous, time-dependent materials, with parameters $\alpha(t)$ and $\beta(t)$. For this situation we cannot use equations (5) and (6). Instead, using equations (3) and (4), we now have ${\bf\nabla}P={\bf\nabla}(\frac{1}{\alpha}U)=\frac{1}{\alpha}{\bf\nabla}U$ and ${\bf\nabla}\cdot{\bf Q}={\bf\nabla}\cdot(\frac{1}{\beta}{\bf V})=\frac{1}{\beta}{\bf\nabla}\cdot{\bf V}$. Hence, elimination of $P$ and ${\bf Q}$ from equations (1) and (2) yields
@@
This system of equations for the wave field quantities $U$ and ${\bf V}$ underlies wave propagation in homogeneous, time-dependent materials. They form the basis for the discussion in the remainder of this paper. For wave propagation in materials that are simultaneously space- and time-dependent, see Innanen (2018); Caloz & Deck-Léger (2020b); Apffel & Fort (2022); Van Manen et al. (2024); Ammari et al. (2024). For non-reciprocal wave propagation in modulated metamaterials, we refer to Lurie (2007); Willis (2011, 2012); Trainiti & Ruzzene (2016); Haberman & Guild (2016); Nassar et al. (2017b, a); Goldsberry et al. (2019); Sotoodehfar et al. (2023).
Equations (7) and (8) are used to derive various fundamental properties of wave propagation and scattering in time-dependent materials. It is also useful to consider a second order wave equation for the scalar field $U({\bf x},t)$. This equation is obtained by eliminating ${\bf V}({\bf x},t)$ from equations (7) and (8), which yields
@@
and ${\bf\nabla}^{2}={\bf\nabla}\cdot{\bf\nabla}=\partial_{x}^{2}+\partial_{z}^{2}$. Finally, we define a time-dependent quantity $\eta(t)$ as
@@ 3.1 Transmission and reflection coefficients
We review transmission and reflection coefficients for a monochromatic plane wave that is incident on a time boundary (Xiao et al., 2014; Morgenthaler, 1958; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b). For convenience we define the time boundary at $t=0$. For $t<0$ we assume constant parameters $\alpha_{1}$ and $\beta_{1}$. Similarly, for $t>0$ we assume constant parameters $\alpha_{2}$ and $\beta_{2}$. For $t<0$, the incident monochromatic plane wave is described in complex notation by
@@
These equations should hold for all ${\bf x}$, which implies ${\bf k}_{1}={\bf k}_{2}$. From this it follows that the angular frequency of the transmitted and reflected waves is different from that of the incident wave, according to $\omega_{2}=\frac{c_{2}}{c_{1}}\omega_{1}$ (Apffel & Fort, 2022; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b). Substituting the appropriate expressions into equations (19) and (20) yields
@@ 3.2 Conservation of net field-momentum density
Next to the net power-flux density, we define the net field-momentum density for electromagnetic waves as $M({\bf x},t)=\Re({\bf D}^{*}\times{\bf B})\cdot{\bf n}$ (Burns et al., 2020; Feynman et al., 1963). For transverse electric and transverse magnetic waves this can be written as
$M({\bf x},t)=\Re(U^{*}{\bf V})\cdot{\bf n}.$ (29)
@@
We analyse the quantities $j({\bf x},t)$ and $M({\bf x},t)$ for monochromatic plane waves before and after a time boundary (equations (13) – (18)). Substituting $U=U_{I}$, ${\bf V}={\bf V}_{I}$ for $t<0$ and $U=U_{T}+U_{R}$, ${\bf V}={\bf V}_{T}+{\bf V}_{R}$ for $t>0$ into equation (29) we find, using equation (25), that $M({\bf x},t)$ is constant for all ${\bf x}$ and $t$. Hence, the net field-momentum density is conserved when passing a time boundary. Using equations (3), (4) and (10), it follows that $j({\bf x},t)$ and $M({\bf x},t)$ are mutually related via $j({\bf x},t)=\frac{c^{2}(t)}{2}M({\bf x},t)$. Hence, $j_{2}=\frac{c_{2}^{2}}{c_{1}^{2}}j_{1}$, where $j_{1}$ and $j_{2}$ are the net power-flux densities before and after the time boundary. Hence, the net power-flux density is not conserved when passing a time boundary. This is the result of energy being added to or extracted from the wave field by the mechanism that modulates the material parameters (Morgenthaler, 1958; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b).
@@ 4 Green’s function
We consider again a homogeneous, time-dependent material, with parameters $\alpha(t)$ and $\beta(t)$. We introduce the Green’s function ${\cal G}({\bf x},{\bf x}_{0},t,t_{0})$ as the response to an impulsive point source at ${\bf x}={\bf x}_{0}$ at time instant $t=t_{0}$. Hence, it obeys wave equation (9), with the source term on the right-hand side replaced by an impulsive point source, according to
@@
Figure 1: Snapshots of the Green’s function ${\cal G}({\bf x},{\bf 0},t,0)$ (convolved with a spatial wavelet) for a piecewise constant time-dependent material, with a time boundary at $t=50.0\,\mu$s. Movie available at https://www.keeswapenaar.nl/TimeMaterial/Green.mp4.
Figure 2: Cross-sections at $x=0$ of the Green’s function of Figure 1, i.e., ${\cal G}({\bf x},{\bf 0},t,0)|_{x=0}$.
We discuss a numerical example of ${\cal G}({\bf x},{\bf x}_{0},t,t_{0})$ for a piecewise constant material with a single time boundary, with propagation velocities $c_{1}=1500$ m/s for $t<50\,\mu$s and $c_{2}=2500$ m/s for $t>50\,\mu$s (the parameter $\beta$ is constant). We choose the source at ${\bf x}_{0}={\bf 0}$ and $t_{0}=0$. Figure 1 shows snapshots of ${\cal G}({\bf x},{\bf 0},t,0)$ (convolved with a 2D spatial wavelet with a central wavenumber $k_{0}/2\pi=200$ m-1). The first three frames show an expanding circular wave front, up to the time boundary at $t=50\,\mu$s. In the following frames the wave field is split into a transmitted wave front, which expands further, and a reflected wave front, which propagates back and collapses to a focus at ${\bf x}={\bf 0}$ and $t=79.75\,\mu$s. At exactly $t=80\,\mu$s the focused field is zero (not shown) and changes its sign, after which it continues as an expanding circular wave front with opposite amplitude, as shown in the last frame. Hence, the focus at ${\bf x}={\bf 0}$ and $t=80\,\mu$s acts as a virtual source for a wave field in a material with propagation velocity $c_{2}=2500$ m/s. Figure 2 shows cross-sections at $x=0$ of these snapshots in the form of a $z,t$-diagram (i.e., for ${\bf x}=(0,z)$, as if receivers were placed along a vertical line in the middle of these snapshots, through the source position). Left of the green line at $t_{0}=0$ the field is zero, conform the causality condition of equation (31). Right of this green line, direct waves propagate away from the source, in two directions (up and down), until they reach the time boundary at $t=50\,\mu$s, indicated by the dashed blue line. Right of this blue line, transmitted and reflected waves can be seen. The reflected waves propagate back and focus at the position of the original source at $z=0$.
@@
Figure 3: The wave field $U({\bf x},t)$ in response to a distributed source in the same piecewise constant time-dependent material as in Figure 1. Movie available at https://www.keeswapenaar.nl/TimeMaterial/TUD.mp4.
Figure 3 is an illustration of equation (34) for a source distribution $s({\bf x},t)=s_{0}({\bf x})\delta(t)$, i.e., for a distributed source $s_{0}$ as a function of ${\bf x}$ at $t=0$. The first frame shows the response at $t=0.25\,\mu$s to this source in the same time-dependent material as in the previous example. The next frames show the evolution of this response over time. After the time boundary at $t=50\,\mu$s, the transmitted field continues expanding, whereas the reflected field propagates back and collapses to a reproduction of the distributed source at $t=79.75\,\mu$s. After this, the focused field changes its sign (at $t=80\,\mu$s) and expands again. Results like this have been obtained with physical experiments by Bacot et al. (2016); Fink & Fort (2017); Peng et al. (2020).
@@
where $H(t)$ is the Heaviside step function. The direct waves in Figure 2 can be seen as an illustration of this expression for ${\cal G}({\bf x},{\bf 0},t,0)|_{x=0}$, convolved with a spatial wavelet.

### v2

@@ 1 Introduction
Temporal changes of material parameters have an effect on wave propagation and scattering, comparable with, but not identical to, the effect of spatial changes. Since the initial work on wave propagation and scattering in time-dependent materials (Morgenthaler, 1958; Jiang, 1975), the research in this field has recently got significant momentum as a result of advances in the engineering of dynamic metamaterials (Caloz & Deck-Léger, 2020a). Most applications deal with electromagnetic waves (Koutserimpas & Fleury, 2018; Mounaix et al., 2020; Ramaccia et al., 2020; Garnier, 2021; Apffel & Fort, 2022; Moussa et al., 2023; Ptitcyn et al., 2023), but mechanical wave propagation and scattering in time-dependent materials is also a rapidly emerging research area. It is proposed as an alternative way of doing ultrasonic time-reversal experiments (Bacot et al., 2016; Fink & Fort, 2017; Bal et al., 2019; Peng et al., 2020; Hidalgo-Caballero et al., 2023) and its potential use in seismic imaging and monitoring is being investigated (Innanen, 2018; Huang et al., 2022).
Various authors discuss the analogy between the wave equations for time-dependent and space-dependent materials (Mendonça & Shukla, 2002; Xiao et al., 2014; de Hoop & Lager, 2014; Salem & Caloz, 2015; Torrent et al., 2018; Caloz & Deck-Léger, 2020b; Van Manen et al., 2024). For 1D applications, the roles of the time- and space-coordinates are interchanged between the wave equations for both types of material (Salem & Caloz, 2015; Torrent et al., 2018; Caloz & Deck-Léger, 2020b; Van Manen et al., 2024). However, causality conditions apply to the time coordinate in both types of material and hence they are not interchanged. As a consequence, solutions of the wave equation in a time-dependent material are, in general, not one-to-one exchangeable with those in a space-dependent material. In a recent paper (Wapenaar et al., 2024) we systematically analyze the similarities and differences of waves in 1D space-dependent and time-dependent materials. In multi-dimensional materials, the number of space dimensions (two or three) is different from the number of time dimensions (one) and hence the analogy between the wave equations for both types of material is less straightforward.
In the current paper we discuss fundamental aspects of wave propagation and scattering in 2D time-dependent materials. The discussion partly reviews work discussed in the references mentioned above but also contains new results. Like in our paper on 1D materials (Wapenaar et al., 2024), we use a unified notation which simultaneously captures electromagnetic, acoustic and elastodynamic shear waves. We discuss transmission and reflection coefficients, Green’s functions, propagation invariants, reciprocity theorems, wave field representations, Green’s function retrieval, propagator matrices and the relation between the Green’s function and the propagator matrix of a time-dependent material.
@@ 2 Unified wave equation
Table 1: Specification of the quantities in equations (1) – (4). For TE (transverse electric) and TM (transverse magnetic) waves, the quantities are electric and magnetic flux densities ${D}$ and ${B}$, electric and magnetic field strengths ${E}$ and ${H}$, permittivity $\varepsilon$, permeability $\mu$, and external electric and magnetic current densities ${J}^{\rm e}$ and ${J}^{\rm m}$. For acoustic waves they are dilatation $\Theta$, mechanical momentum density ${m}$, acoustic pressure $p$, particle velocity ${v}$, compressibility $\kappa$, mass density $\rho$, volume-injection rate density $q$ and external force density ${f}$. For SH (horizontally polarised shear) waves, $m$, $v$, $\rho$ and $f$ are defined the same as for acoustic waves, and the additional quantities are strain $e$, stress $\tau$, shear modulus $\mu$ and external deformation rate density $h$.
@@ 3. Acoustic $-\Theta$ $m_{x}$ $m_{z}$ $p$ $v_{x}$ $v_{z}$ $\kappa$ $\rho$ $q$ $f_{x}$ $f_{z}$
where $\alpha$ and $\beta$ are material parameters. They are also specified in Table 1 for the different wave phenomena. We can use these equations to eliminate two of the four wave field quantities from equations (1) and (2). We consider three cases.
  1. 1.
    Inhomogeneous, time-dependent material, with parameters $\alpha({\bf x},t)$ and $\beta({\bf x},t)$.
    Substituting equations (3) and (4) straightforwardly into equations (1) and (2) yields
    $\partial_{t}(\alpha P)+{\bf\nabla}\cdot{\bf Q}$ $=$ $a,$ (5)
    $\partial_{t}(\beta{\bf Q})+{\bf\nabla}P$ $=$ ${\bf b}.$ (6)
    For example, for the special situation of acoustic waves (row 3 in Table 1) this gives $\partial_{t}(\kappa p)+{\bf\nabla}\cdot{\bf v}=q$ and $\partial_{t}(\rho{\bf v})+{\bf\nabla}p={\bf f}$ (Nassar et al., 2017b).
  2. 2.
    Inhomogeneous, time-independent material, with parameters $\alpha({\bf x})$ and $\beta({\bf x})$.
    For this situation we have $\partial_{t}U=\partial_{t}(\alpha P)=\alpha\partial_{t}P$ and $\partial_{t}{\bf V}=\partial_{t}(\beta{\bf Q})=\beta\partial_{t}{\bf Q}$. Hence, we obtain from equations (5) and (6)
@@
    As an example, for acoustic waves these expressions become $\kappa\partial_{t}p+{\bf\nabla}\cdot{\bf v}=q$ and $\rho\partial_{t}{\bf v}+{\bf\nabla}p={\bf f}$.
    The well-known system of equations (7) and (8) for the wave field quantities $P$ and ${\bf Q}$ underlies wave propagation in inhomogeneous, time-independent materials with parameters $\alpha({\bf x})$ and $\beta({\bf x})$, which may vary continuously with space. When the material contains space boundaries with normal ${\bf n}({\bf x})$, equations (7) and (8) are supplemented with boundary conditions. The boundary conditions state that $P$ and ${\bf Q}\cdot{\bf n}$ are continuous over those space boundaries.
  3. 3.
    Homogeneous, time-dependent material, with parameters $\alpha(t)$ and $\beta(t)$.
    For this situation we cannot use equations (7) and (8). Instead, using equations (3) and (4), we now have ${\bf\nabla}P={\bf\nabla}(\frac{1}{\alpha}U)=\frac{1}{\alpha}{\bf\nabla}U$ and ${\bf\nabla}\cdot{\bf Q}={\bf\nabla}\cdot(\frac{1}{\beta}{\bf V})=\frac{1}{\beta}{\bf\nabla}\cdot{\bf V}$. Hence, elimination of $P$ and ${\bf Q}$ from equations (1) and (2) yields
@@
    For example, for acoustic waves these expressions read $-\partial_{t}\Theta+\frac{1}{\rho}{\bf\nabla}\cdot{\bf m}=q$ and $\partial_{t}{\bf m}-\frac{1}{\kappa}{\bf\nabla}\Theta={\bf f}$.
    The system of equations (9) and (10) for the wave field quantities $U$ and ${\bf V}$ underlies wave propagation in homogeneous, time-dependent materials with parameters $\alpha(t)$ and $\beta(t)$, which may vary continuously with time. When the material contains time boundaries, equations (9) and (10) are supplemented with boundary conditions. The boundary conditions state that $U$ and ${\bf V}$ are continuous over those time boundaries (Xiao et al., 2014; Morgenthaler, 1958; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b; de Hoop & Lager, 2014).
Although equations (5) and (6) hold for materials that are simultaneously space- and time-dependent, we restrict the discussion in this paper to homogeneous time-dependent materials, governed by equations (9) and (10), supplemented with boundary conditions for $U$ and ${\bf V}$. This allows the translation of several fundamental concepts known for waves in inhomogeneous time-independent materials (case ii) to waves in homogeneous time-dependent materials (case iii). For aspects of wave propagation in materials that are simultaneously space- and time-dependent (case i), see references Caloz & Deck-Léger (2020b); Apffel & Fort (2022); Van Manen et al. (2024); Ammari et al. (2024). For non-reciprocal wave propagation in metamaterials, we refer to Lurie (2007); Willis (2011, 2012); Trainiti & Ruzzene (2016); Haberman & Guild (2016); Nassar et al. (2017b, a); Goldsberry et al. (2019); Sotoodehfar et al. (2023).
Next to the first-order equations (9) and (10), it is also useful to consider a second order wave equation for the scalar field $U({\bf x},t)$. This equation is obtained by eliminating ${\bf V}({\bf x},t)$ from equations (9) and (10), which yields
@@
and ${\bf\nabla}^{2}={\bf\nabla}\cdot{\bf\nabla}=\partial_{x}^{2}+\partial_{z}^{2}$. Note that, unlike in the wave equation for an inhomogeneous time-independent material, in which a material parameter appears between the space-derivatives, in equation (11) a material parameter appears between the time-derivatives (Zurita-Sánchez et al., 2009; Torrent et al., 2018). Finally, we define a time-dependent quantity $\eta(t)$ as
@@ 3.1 Transmission and reflection coefficients
We review transmission and reflection coefficients for a monochromatic plane wave that is incident on a time boundary (Xiao et al., 2014; Morgenthaler, 1958; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b). For convenience we define the time boundary at $t=0$. For this analysis, the medium parameters $\alpha(t)$ and $\beta(t)$ are step functions of time, with a finite jump at $t=0$. For $t<0$ we have time-independent parameters $\alpha_{1}$ and $\beta_{1}$. Similarly, for $t>0$ we have time-independent parameters $\alpha_{2}$ and $\beta_{2}$. For $t<0$, the incident monochromatic plane wave is described in complex notation by
@@
These equations should hold for all ${\bf x}$, which implies ${\bf k}_{1}={\bf k}_{2}$. From this it follows that the angular frequency of the transmitted and reflected waves is different from that of the incident wave, according to $\omega_{2}=\frac{c_{2}}{c_{1}}\omega_{1}$ (Apffel & Fort, 2022; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b). In other words, frequency conversion occurs at a time boundary. Substituting the appropriate expressions into equations (21) and (22) yields
@@ 3.2 Conservation of net field-momentum density
Next to the net power-flux density, we define the net field-momentum density for electromagnetic waves as $M({\bf x},t)=\frac{1}{2}\Re({\bf D}^{*}\times{\bf B})\cdot{\bf n}$ (Burns et al., 2020; Feynman et al., 1963). For transverse electric and transverse magnetic waves this can be written as
$M({\bf x},t)=\frac{1}{2}\Re(U^{*}{\bf V})\cdot{\bf n}.$ (31)
@@
We analyse the quantities $j({\bf x},t)$ and $M({\bf x},t)$ for monochromatic plane waves before and after a time boundary (equations (15) – (20)). Substituting $U=U_{I}$, ${\bf V}={\bf V}_{I}$ for $t<0$ and $U=U_{T}+U_{R}$, ${\bf V}={\bf V}_{T}+{\bf V}_{R}$ for $t>0$ into equation (31) we find, using equation (27), that $M({\bf x},t)$ is constant for all ${\bf x}$ and $t$. Hence, the net field-momentum density is conserved when passing a time boundary. Using equations (3), (4) and (12), it follows that $j({\bf x},t)$ and $M({\bf x},t)$ are mutually related via $j({\bf x},t)=c^{2}(t)M({\bf x},t)$. Hence, $j_{2}=\frac{c_{2}^{2}}{c_{1}^{2}}j_{1}$, where $j_{1}$ and $j_{2}$ are the net power-flux densities before and after the time boundary. Hence, the net power-flux density is not conserved when passing a time boundary. This is the result of energy being added to or extracted from the wave field by the mechanism that modulates the material parameters (Morgenthaler, 1958; Mendonça & Shukla, 2002; Caloz & Deck-Léger, 2020b).
@@ 4 Green’s function
We consider a homogeneous, time-dependent material, with arbitrary piecewise continuous parameters $\alpha(t)$ and $\beta(t)$ (case iii in section 2). We introduce the Green’s function ${\cal G}({\bf x},{\bf x}_{0},t,t_{0})$ as the response to an impulsive point source at ${\bf x}={\bf x}_{0}$ at time instant $t=t_{0}$, observed at position ${\bf x}$ and time $t$. Hence, it obeys wave equation (11), with the source term on the right-hand side replaced by an impulsive point source, according to
@@
Figure 1: Snapshots of the Green’s function ${\cal G}({\bf x},{\bf 0},t,0)$ (convolved with a spatial wavelet) for a piecewise constant time-dependent material, with a time boundary at $t=50.0\,\mu$s (see also movie Green.mp4 in the online material).
Figure 2: (a) Cross-sections at $x=0$ of the Green’s function of Figure 1, i.e., ${\cal G}({\bf x},{\bf 0},t,0)|_{x=0}$. (b) Scaled amplitude cross-sections of (a), measured along the direct and transmitted waves (blue) and along the reflected wave (red).
We discuss a numerical example of ${\cal G}({\bf x},{\bf x}_{0},t,t_{0})$ for a piecewise constant material with a single time boundary at $t=50\,\mu$s, with propagation velocities $c_{1}=1500$ m/s for $t<50\,\mu$s and $c_{2}=2500$ m/s for $t>50\,\mu$s (the parameter $\beta$ is constant). The transmission and reflection coefficients at the time boundary, as defined in equations (25) and (26), are $T_{u}=0.8$ and $R_{u}=0.2$, respectively. We choose the source at ${\bf x}_{0}={\bf 0}$ and $t_{0}=0$. Figure 1 shows snapshots of ${\cal G}({\bf x},{\bf 0},t,0)$ (convolved with a 2D spatial wavelet with a central wavenumber $k_{0}/2\pi=100$ m-1). The first three frames show an expanding circular wave front, up to the time boundary at $t=50\,\mu$s. In the following frames the wave field is split into a transmitted wave front, which expands further, and a reflected wave front, which propagates back and collapses to a focus at ${\bf x}={\bf 0}$ and $t=79.75\,\mu$s. At exactly $t=80\,\mu$s the focused field is zero (not shown) and changes its sign, after which it continues as an expanding circular wave front with opposite amplitude, as shown in the last frame. Hence, the focus at ${\bf x}={\bf 0}$ and $t=80\,\mu$s acts as a virtual source for a wave field in a material with propagation velocity $c_{2}=2500$ m/s. Figure 2(a) shows cross-sections at $x=0$ of these snapshots in the form of a $z,t$-diagram (i.e., for ${\bf x}=(0,z)$, as if receivers were placed along a vertical line in the middle of these snapshots, through the source position). Left of the green line at $t_{0}=0$ the field is zero, conform the causality condition of equation (33). Right of this green line, direct waves propagate away from the source, in two directions (up and down), until they reach the time boundary at $t=50\,\mu$s, indicated by the dashed blue line. Right of this blue line, transmitted and reflected waves can be seen. Tracing the transmitted waves back in time along the purple dashed rays, they appear to originate from a virtual source at $z=0$ and $t_{V}=20\,\mu$s. The reflected waves propagate back and focus at the position of the original source at $z=0$ and $t_{F}=80\,\mu$s. This focus acts as a virtual source for the reflected waves beyond $t_{F}=80\,\mu$s. Figure 2(b) shows an amplitude cross-section of Figure 2(a), scaled such that the amplitude of the direct wave arriving at the time boundary at $t=50\,\mu$s equals 1. The blue curve shows that the amplitudes of the direct and transmitted waves gradually decay with time, and jump from 1 to $T_{u}=0.8$ across the time boundary. The red curve shows that the reflected wave starts directly after the time boundary with an amplitude of $R_{u}=0.2$. Its amplitude initially increases during back-propagation, changes sign at the focus and decreases after the focus.
@@
Figure 3: The wave field $U({\bf x},t)$ in response to a distributed source in the same piecewise constant time-dependent material as in Figure 1 (see also movie TUD.mp4 in the online material).
Figure 3 is an illustration of equation (36) for a source distribution $s({\bf x},t)=s_{0}({\bf x})\delta(t)$, i.e., for a distributed source $s_{0}$ as a function of ${\bf x}$ at $t=0$ (this time the 2D spatial wavelet has a central wavenumber $k_{0}/2\pi=200$ m-1). The first frame shows the response at $t=0.25\,\mu$s to this distributed source in the same time-dependent material as in the previous example. The next frames show the evolution of this response over time. After the time boundary at $t=50\,\mu$s, the transmitted field continues expanding, whereas the reflected field propagates back and collapses to a reproduction of the distributed source at $t=79.75\,\mu$s. After this, the focused field changes its sign (at $t=80\,\mu$s) and expands again. Results like this have been obtained with physical experiments by Bacot et al. (2016); Fink & Fort (2017); Peng et al. (2020).
@@
where $H(t)$ is the Heaviside step function. The direct waves in Figure 2(a) can be seen as an illustration of this expression for ${\cal G}({\bf x},{\bf 0},t,0)|_{x=0}$, convolved with a spatial wavelet.